Inside a lift accelerating upwards at 2m/s^2 a loose bolt falls from the ceiling. If height of lift is 2m. Find the speed of bolt with which it hits the floor. Initial speed of lift is 5m/s in upward direction.In this question i found the answer as velocity of bolt wrt lift as 4(1/3)^1/2 m/s.My doubt is what will be the velocity of bolt wrt ground ? The acceleration of bolt wrt ground and why ......plz explain

Grade:11

1 Answers

Vikas TU
7 years ago
From the Relative velocity and accln. concept,
assign the velocities and acclns. to the bolt in opposite direction of the lift and apply  the Newton’s aw of eqns.
Therefore,
v^2 = (0-5)^2 + 2(g + 2)*2
v = root(73) => 856 m/s.
accln. = (a+g) => 12 m/s^2